OpenStudy (anonymous):

Find the 90 th derivative of the function ƒ(x)=cos(x) The answer is function?

OpenStudy (anonymous):

every 4th derivative is cos x 90= 2(mod 4).. so, 90th derivative will be 2nd derivative of cos x, i.e.. -cos x

OpenStudy (anonymous):

Thanks!!!

OpenStudy (experimentx):

f1(x) = - sin(x) f2(x) = -cos(x) f3(x) = sin(x) f4(x) = cos(x) .. .. f4n(x) = cos(X) f4n+1(x) = sin(X) f4n-2(x) = -cos(X) f4n-1(x) = -sin(X) I don't seem to find the general governing rule for all